Can someone help me on this problem?
I need to solve the following equation for x, and show proof:
2(x-3) - 3(x + 5) = 3(x-2) - 7
Thanks,
Sher

Answer by ankor@dixie-net.com(22740)

2(x-3) - 3(x + 5) = 3(x-2) - 7
:
Mult what's inside the brackets, remember to change the sign inside the brackets if there is a negative outside the brackets
:
2x - 6 - 3x - 15 = 3x - 6 - 7
:
generally we perform operations to get the unknown (x) on the left and the numerical values on the right. (Altho it's Ok to reverse that if you like)
:
2x - 3x - 3x = -6 -7 + 6 + 15
:
-4x = +8
:
x = +8/-4
:
x = -2
:
You can prove your solution by substituting -2 for x in the original equation and see if equality reigns
:
2(x-3) - 3(x + 5) = 3(x-2) - 7
:
2(-2 -3) - 3(-2 +5) = 3(-2-2) - 7
:
2(-5) - 3(+3) = 3(-4) - 7; did the math inside the brackets
:
-10 - 9 = - 12 - 7
:
-19 = -19 proves that x = -2